OpenStudy (anonymous):

If anyone knows anything about chemistry do you think you could help me? How do you find the number of neutrons in an atom?

OpenStudy (anonymous):

An atom is described by two different numbers. the first is known as the atomic number. The second is known as the or atomic mass. The atomic number will be the number of protons in an atom, and determines what that atom is. The atomic mass however is the number of nucleons within an atom. A nucleon is just a collective name for protons and neutrons. So the number of neutrons in any atom will be the difference between the Atomic Mass and the Atomic number. For example an isotope of iron has an atomic number of 26, but an atomic mass of 56. Hence there will be 26 protons in the nucleus, and 56-26=30 neutrons.

OpenStudy (anonymous):

Subtract the atomic number from mass number and you get number of neutrons.Atomic number is always lower than mass number so lets consider an element of atomic number=26 and mass number=85. Number of neutrons(n)=Mass Number(A)- Atomic Number(Z) n=85-26 n=59 Number of neutrons=59 We are subtracting number of electrons and protons from the sum of protons/electrons and neutrons.If you will subtract electrons/protons from the mass number i.e the sum of electron/proton and neutrons then you will obviously get total number of neutrons.